# Quantifying Fire Risk Index in Chemical Industry Using Statistical Modeling Procedure

### v1

@@ Abstract
Fire incident reports contain detailed textual narratives that capture causal factors often overlooked in structured records, while financial damage amounts provide measurable outcomes of these events. Integrating these two sources of information is essential for uncovering interpretable links between descriptive causes and their economic consequences. To this end, we develop a data-driven framework that constructs a composite Risk Index, enabling systematic quantification of how specific keywords relate to property damage amounts. This index facilitates both the identification of high-impact terms and the aggregation of risks across semantically related clusters, thereby offering a principled measure of fire-related financial risk. Using more than a decade of Korean fire investigation reports on the chemical industry classified as Special Buildings (2013–2024), we employ topic modeling and network-based embedding to estimate semantic similarities from interactions among words, and subsequently apply Lasso regression to quantify their associations with property damage amounts, thereby estimate fire risk index. This approach enables us to assess fire risk not only at the level of individual terms but also within their broader textual context, where highly interactive related words provide insights into collective patterns of hazard representation and their potential impact on expected losses. The analysis highlights several domains of risk, including hazardous chemical leakage, unsafe storage practices, equipment and facility malfunctions, and environmentally induced ignition. The results demonstrate that text-derived indices provide interpretable and practically relevant insights, bridging unstructured narratives with structured loss information and offering a basis for evidence-based fire risk assessment and management.
@@ 1 Introduction
Fire accidents pose significant threats to human life, property, and industrial activities. In particular, facilities in the chemical industry are highly vulnerable due to the presence of flammable and hazardous materials, which can lead to catastrophic outcomes once a fire occurs. Understanding the underlying risk factors of such accidents is therefore crucial for developing effective prevention strategies, designing safety regulations, and minimizing social and economic damages.
While structured fire statistics and engineering-based indicators provide valuable information, they often fail to capture the nuanced circumstances surrounding each incident. In contrast, post-incident investigation records written by frontline fire officers contain rich textual descriptions of ignition causes, equipment failures, material involvement, and contextual human or environmental factors observed on site. These narratives represent first-hand accounts of accident circumstances, preserving details that are easily overlooked or lost when data are reduced to predefined categories. Moreover, examining textual data reveals co-occurrence patterns among words, indicating that fire-related factors often appear in dependent and interrelated structures. Such contextual associations can only be identified through systematic analysis of unstructured narratives, enabling a deeper understanding of how multiple risk factors interact to influence the likelihood and severity of fire accidents. Analyzing such textual evidence is therefore essential not only for uncovering hidden or emerging risk factors, but also for linking causal mechanisms with the resulting scale of damage.
@@
Previous research has employed a variety of approaches to analyze fire-related risks. Text-based studies have demonstrated the potential of mining unstructured documents. For instance, Kim and Hwang [10] applied topic modeling techniques to accident verdicts from ship fire cases to identify ignition sources, flammable materials, and negligence-related causes. While this study represented an important step in applying text analytics to fire accidents, it primarily relied on word frequency clustering and did not capture the correlated structure among words that reflects the interdependent nature of fire risk factors. Similarly, Tirunagari [20] investigated maritime accident investigation reports using text mining to extract causal relations among contributing factors. Although not focused on fires, this study highlights the methodological potential of analyzing free-form investigation records, underscoring the value of unstructured narratives in uncovering complex causal structures in accident data.
@@
Structured and indicator-based approaches have also received considerable attention. Ma et al. [14] conducted a comprehensive data-driven analysis of over one million building fire reports, integrating structured incident records with socioeconomic and structural variables to assess the effects of detection systems and extinguishing devices on fire spread and injury risk. Although similar in its goal of combining textual and structured sources, its textual component was limited to predefined categorical fields rather than free-text narratives, preventing it from capturing word co-occurrence patterns or the indexical role of terms. In a related effort, Zhang et al. [25] developed an indicator system for urban fire risk assessment, emphasizing meteorological conditions and building characteristics. Yet, this approach excluded unstructured textual data, thereby omitting rich qualitative details from fire incident reports.
Finally, research has also explored narrative-based perspectives on fire events. Russo et al. [17] analyzed wildfire narratives to identify and characterize multiple social storylines concerning causes, consequences, and potential solutions. This work illustrates how unstructured narratives can provide a richer and more contextualized understanding of fire events by linking causal descriptions to broader societal implications.
Taken together, these studies highlight the promise of both textual and structured approaches for understanding fire risks, but also reveal important limitations. Existing work has seldom analyzed fire investigation narratives written directly by frontline fire officers, which contain detailed accounts of ignition mechanisms, equipment failures, and contextual factors surrounding each incident. To address this gap, our study develops a framework that systematically analyzes such unstructured text while explicitly accounting for the dependent structure among words, thereby identifying coherent topics that represent interrelated fire causes. Furthermore, by integrating this textual analysis with structured data on fire property losses, we construct a risk index factor that quantifies the degree of fire risk across different categories of incidents. This combined approach not only facilitates the extraction of meaningful risk factors through salient keywords, but also provides researchers and practitioners with an interpretable index that reflects both causal conditions and damage outcomes. Ultimately, the proposed framework enables a more comprehensive and data-driven understanding of fire risk, helping stakeholders identify and prioritize critical factors for prevention and mitigation.
@@ 2 Materials and Methods
In the third step (Step 3), the latent item response model [8] is applied to infer the positions of words in a continuous latent space, which captures their semantic relationships. The estimated distances between words can be interpreted as measures of semantic similarity: for example, if two words are placed close to each other in the latent space, they are more likely to represent semantically related concepts. Leveraging these latent positions, words are clustered into semantically coherent groups, thereby enabling the construction of interpretable clusters of risk-related vocabulary.
@@ 2.1 Data
The dataset used in this study consists of records from post-incident investigations of fire accidents in these chemical industry Special Buildings between 2013 and 2024. The fire investigation records are unstructured text documents describing the fire circumstances, such as the cause, involved equipment, and ignited materials. All investigation texts are written in Korean. In addition, the dataset includes property damage estimates provided by local fire departments, covering both movable and immovable property losses.
@@ 4. Step 4
The resulting joint likelihood over all biterms $B$ is:
@@
The conditional posterior for topic assignment $z$ is:
@@
where $n_{z}$ is the number of biterms currently assigned to topic $z$, $n_{w|z}$ is the number of times word $w$ is assigned to topic $z$, and $\mathbf{z}_{-b}$ denotes topic assignments excluding the current biterm. To address convergence issues that may arise from direct Gibbs sampling, we employ collapsed Gibbs sampling, which integrates out $\bm{\theta}$ and $\bm{\phi}_{z}$ by exploiting conjugacy of the Dirichlet–Multinomial distributions [13]. After sufficient sampling iterations, the posterior estimates of topic–word distributions and topic proportions are:
@@
After inference, the BTM yields both $\bm{\phi}_{w|z}$ and $\bm{\theta}_{z}$. Each topic is defined by its associated words and their probabilities, making it possible to compare topics by identifying their distinguishing word distributions. Because many words may appear across multiple topics with non-negligible probabilities, we focus on representative terms. Specifically, we construct a word–topic probability matrix $\mathbf{X}$ of dimension $N\times P$, where $N$ is the number of words and $P$ is the number of topics. To identify characteristic words, we compute two measures for each row of $\mathbf{X}$: (i) the coefficient of variation across topics, and (ii) the maximum probability. The coefficient of variation captures the relative dispersion of a word’s probabilities across topics, while the maximum probability identifies whether the word is strongly associated with at least one topic. Words with both high dispersion and high maximum probability are retained as representative terms, enabling more interpretable characterization of latent topics.
@@ 2.3 Clustering Topics via Latent Space Item Response Model
where $x_{i,p}$ denotes the probability of word $i$ belonging to topic $p$, $\bm{a}_{i}$ and $\bm{b}_{p}$ are attribute parameters, and ${\bf v}_{i}-{\bf u}_{p}||$ captures the Euclidean distance between the latent positions of topic $p$ and word $i$. A shorter distance implies stronger association between the word and the topic. Bayesian inference is employed to estimate the full parameter set $\bm{\Theta}=\{\bm{a},\bm{b},\mathbf{U},\mathbf{V}\}$ with appropriate priors, and parameters are sampled using Markov chain Monte Carlo (MCMC). As a result, we obtain latent positions of topics $\mathbf{v}_{i}$ in $\mathbb{R}^{d}$, forming the topic coordinate matrix ${\bf A}\in\mathbb{R}^{d\times P}$.
@@ 3.1 Topic Characterization through Words
Table 1 summarizes the thematic interpretation of topics extracted from the fire investigation records. In addition, Table 2 presents the distribution of words, listing the top five terms with the highest probabilities for each topic. These high-probability keywords serve as representative indicators of topic relevance and provide the basis for characterizing the underlying thematic structure.
@@
Table 2: Top five representative words for each extracted topic based on topic–word probability distributions. These high-probability words provide the basis for interpreting the thematic characteristics of the topics.
Topic Top 5 Words Topic 1 cleaning, cleanig room, wastewater treatment plant, agitator, machine room Topic 2 mixer, drum can, pallet, plastic, explosion Topic 3 electrical short-circuit, electrical distribution, flame, circuit breaker, SWGR Topic 4 extractor hood, oven, laboratory, transformer, electricity Topic 5 heating wire, outdoor unit of air-conditioner, thermal/acoustical insulation, air-conditioner, rooftop Topic 6 dust collection equipment, duct, flame, plenty of, filter system Topic 7 storage, sesame dregs, residues, spontaneous combustion, storage room Topic 8 forklift, battery, electrical short-circuit, distribution, engine compartment Topic 9 explosion, container, reactor, oil vapor, static electricity Topic 10 dust collection equipment, duct, flame, ventilation duct, plastic Topic 11 drying equipment, powder, silicon, base material, storage Topic 12 base material, mixing equipment, banbury, thermal oil, flame Topic 13 welding, cooling tower, flame, sandwich panel, acetylene Topic 14 air compressor, motor, electrical short-circuit, vulcanizer, solenoid valve Topic 15 waste materials, cigarette butts, waste, stacking, flame
@@ 3.2 Thematic Aggregation of Topics
Building upon the topic–word distributions, we further inferred interactions among words to explore higher-level thematic structures. Figure 2 illustrates the latent positions of words projected onto a two-dimensional space, where clustering was performed using the $k$-means algorithm. The visualization highlights clusters by distinct colors, showing how semantically related words are grouped in close proximity. Based on model selection criteria, a total of 15 clusters were identified as valid. As shown in the figure, words located near one another in the latent space tend to form coherent clusters, thereby capturing meaningful associations beyond the topic-level representation.
@@
3.3 Cluster Interpretation
According to Table 3, the top 10 keywords in each cluster are ordered by their proximity to other words, as determined by distance calculations based on the estimated latent positions. This ordering enables a more precise characterization of the distinctive features of each cluster. The identified clusters can be grouped into several overarching thematic categories of fire risks.
Chemical and Reaction-Related Risks.
Clusters involving hazardous chemical processes highlight the potential for fires and explosions arising from chemical leaks, abnormal reactions, or spontaneous combustion of unstable substances. For instance, distillation column, exposure, and solvent (Cluster 1) point to explosions due to hazardous material leakage and pressure control failures. Similarly, abnormal reaction terms such as pharmaceutical, film, and epoxy (Cluster 5) suggest severe accidents during abnormal chemical handling. Risks of spontaneous combustion are also evident, as indicated by heat of reaction, expired reagents, and cooking oil (Cluster 4), as well as residues like compost and sesame dregs (Cluster 14), which underscore dangers from abandoned chemicals and oils.
Industrial Work and Machinery-Related Risks.
Another group of clusters reflects ignition sources associated with industrial work environments and machinery. Terms such as polishing, painting, and waste (Cluster 2) emphasize ignition from heat accumulation near combustible interior materials. The presence of incinerate, waste wood, and dust collection equipment (Clusters 3 and 15) indicates residual heat or sparks generated during work with combustibles. Likewise, coating, cutting oil, and grinders (Cluster 7) highlight ignition risks from friction heat or sparks produced by machinery.
Electrical and Equipment-Related Risks.
Electrical fire hazards are represented by clusters centered on terms such as storage, cable, and control box (Cluster 6), which indicate electrical fires in storage or handling areas. Other clusters highlight process-related electrical causes (Cluster 11) and fires occurring during maintenance of machinery (Cluster 12). These categories capture risks arising not from chemical reactions but from equipment and power systems in industrial sites.
Vapor, High-Temperature, and Dust Accumulation Risks.
A further set of clusters concerns ignition in environments with flammable vapors, high temperatures, or combustible dust. For example, vapor, hazardous materials, and grease (Cluster 8) illustrate ignition hazards during the use of organic solvents. Likewise, drying room, cosmetic, and ventilation fan (Cluster 9) point to fire risks in areas with rising oil vapor temperatures. Clusters including dilution, melting fusion, and vulcanizer (Cluster 13) capture ignition due to accumulated combustible dust and oil vapors, further extending the scope of hazards beyond purely chemical origins.
Accumulated Combustibles.
Finally, some risks are associated with the buildup of general combustible materials, as indicated by Cluster 10, which represents fire outbreaks triggered by ignition in accumulated combustibles. This category is particularly relevant to non-chemical industrial environments where storage and waste management play a central role.
@@ 3.3 Cluster Interpretation
Table 3: Clusters of 10 representative keywords derived from inter-word correlations based on estimated latent positions.

### v2

@@ Abstract
Fire incident reports contain detailed textual narratives that capture causal factors often overlooked in structured records, while financial damage amounts provide measurable outcomes of these events. Integrating these two sources of information is essential for uncovering interpretable links between descriptive causes and their economic consequences. To this end, we develop a data-driven framework that constructs a composite Risk Index, enabling systematic quantification of how specific keywords relate to property damage amounts. This index facilitates both the identification of high-impact terms and the aggregation of risks across semantically related clusters, thereby offering a principled measure of fire-related financial risk. Using more than a decade of Korean fire investigation reports on the chemical industry classified as Special Buildings (2013–2024), we employ topic modeling and network-based embedding to estimate semantic similarities from interactions among words, and subsequently apply Lasso regression to quantify their associations with property damage amounts, thereby estimate fire risk index. This approach enables us to assess fire risk not only at the level of individual terms but also within their broader textual context, where highly interactive related words provide insights into collective patterns of hazard representation and their potential impact on expected losses. The analysis highlights several domains of risk, including hazardous chemical leakage, unsafe storage practices, equipment and facility malfunctions, and environmentally induced ignition. The results demonstrate that text-derived indices provide interpretable and practically relevant insights, bridging unstructured narratives with structured loss information and offering a basis for evidence-based fire risk assessment and management. The derived Risk Index provides practical reference data for both safety management and insurance underwriting by enabling the prioritization of preventive measures within industrial sites and offering quantitative guidance for assessing facility-specific risk levels in insurance decisions. An R implementation of the proposed framework is openly available for public use.
@@ 1 Introduction
Fire accidents pose serious threats to industrial safety, human life, and the economy. In particular, chemical facilities are highly vulnerable because of their dependence on flammable and reactive materials, where minor ignition can escalate into catastrophic events. Understanding and quantifying such risks are therefore essential for developing effective preventive measures and for guiding insurance and safety management policies. While structured fire statistics provide valuable information about fire frequency and damages, they often fail to capture the complex causal mechanisms recorded in textual fire investigation narratives. These reports, written by on-site officers, contain unstructured yet highly informative descriptions that reflect contextual factors such as process operations, equipment malfunctions, material leakage, and human or environmental interactions. Leveraging such unstructured narratives through text mining offers a promising pathway for discovering latent risk patterns that cannot be detected from aggregated statistics alone.
@@
Previous research has employed a variety of approaches to analyze fire-related risks. Text-based studies have demonstrated the potential of mining unstructured documents. For instance, Kim and Hwang [16] applied topic modeling techniques to accident verdicts from ship fire cases to identify ignition sources, flammable materials, and negligence-related causes, while Tirunagari [34] investigated maritime accident investigation reports using text mining to extract causal relations among contributing factors. Although these studies represented important early steps in applying text analytics to fire accidents, they primarily relied on word-frequency clustering and did not capture the correlated structure among words that reflects the interdependent nature of fire risk factors. More recently, Liu et al. [24] analyzed scientific literature to trace evolving research trends in forest fire studies, and Zhao et al. [44] constructed a comprehensive database of wildland–urban interface fires. Similar text mining frameworks have been used to identify risk factors in ship fires [43], heritage building fires [36], and industrial settings such as mining and chemical enterprises [37, 22, 25, 41, 5]. Narrative-based perspectives have also been explored: for example, Russo et al. [31] analyzed wildfire narratives to identify multiple social storylines concerning causes, consequences, and potential solutions, illustrating how unstructured narratives can provide a contextualized understanding of fire events and their socio-environmental implications.
@@
Structured and indicator-based approaches have likewise received considerable attention. Ma et al. [26] conducted a data-driven analysis of over one million building fire reports, integrating structured incident records with socioeconomic and structural variables to assess the effects of detection systems and extinguishing devices on fire spread and injury risk. Similarly, Zhang et al. [42] developed an indicator system for urban fire risk assessment that emphasizes meteorological and building characteristics. Other domain-specific efforts include quantitative frameworks for Value-at-Risk of heritage structures [36], climate-related hazard assessment [7], and chemical industry safety evaluation [18]. However, most of these structured approaches exclude unstructured textual data, thereby omitting rich qualitative details that often contain the causal reasoning behind fire events.
Parallel progress in natural language processing and machine learning has broadened the methodological basis for text-based risk analytics. Transformer-based models such as Sentence-BERT have improved short-text clustering by modeling contextual similarity between sentences [17, 28], while hybrid topic–autoencoder architectures enhance detection of latent thematic structures [6, 20]. Graph-based and network-analytic frameworks have also been proposed to represent relationships among semantic entities, enabling discovery of interaction mechanisms in multilingual or domain-specific corpora [12, 15, 45]. These methodological advances have improved text clustering, anomaly detection, and classification across diverse fields, including risk disclosure analysis [2], software fault diagnosis [46], and insurance claim interpretation [29]. Yet, existing fire-related applications remain largely descriptive, focusing on frequent keywords or co-occurrence patterns rather than connecting textual semantics to measurable outcomes such as financial or property loss.
Taken together, the literature reveals a persistent methodological and practical gap. Existing fire text-mining studies seldom model the dependent structure among risk-related terms that jointly describe operational and environmental conditions within accident narratives. Likewise, few studies link textual features to structured indicators such as damage costs or insurance data. Consequently, current approaches cannot quantify how textual expressions of risk translate into tangible loss or prioritize critical risk factors by their financial relevance. Addressing this gap, the present study introduces a unified framework that systematically analyzes fire investigation narratives, estimates semantic dependencies among risk-related terms, and integrates these text-derived features with estimated damages data to construct a composite, loss-aware fire Risk Index. This approach bridges unstructured and structured domains, offering a scalable and interpretable foundation for quantitative fire risk analysis and decision-making.
This framework is expected to be integrated into industrial safety management and insurance risk-assessment systems. The derived Risk Index provides quantitative reference data that enable facility managers to prioritize preventive measures according to site-specific hazards extracted from textual records. It also assists insurers in evaluating facility-level risks that cannot be captured by conventional loss-ratio statistics. By linking narrative investigation data with measurable financial outcomes, the framework facilitates data-driven decision making for both workplace safety management and insurance underwriting.
@@ 2 Materials and Methods
In the third step (Step 3), the latent item response model [13] is applied to infer the positions of words in a continuous latent space, which captures their semantic relationships. The estimated distances between words can be interpreted as measures of semantic similarity: for example, if two words are placed close to each other in the latent space, they are more likely to represent semantically related concepts. Leveraging these latent positions, words are clustered into semantically coherent groups, thereby enabling the construction of interpretable clusters of risk-related vocabulary. Importantly, the estimation of interword interactions accounts for both the frequency and contextual sparsity of co-occurrences. Thus, words are not measured as strongly associated merely because they occasionally appear together; rather, their latent positions are inferred based on consistent and dense interaction patterns with other words across the corpus.
@@ 2.1 Data
The dataset used in this study consists of records from post-incident investigations of fire accidents in these chemical industry Special Buildings between 2013 and 2024. As shown in Table 1, the fire investigation records are unstructured text documents describing the fire circumstances, such as the cause, involved equipment, and ignited materials. All investigation texts are written in Korean. In addition, the dataset includes property damage estimates verified by local fire departments in accordance with the Fire Damage Assessment Manual of the National Fire Agency of Korea, covering both movable and immovable property losses such as buildings, facilities, machinery, and inventory assets.
Table 1: Ten randomly selected representative excerpts of fire investigation records
ID fire investigation records 1 A fire occurred in the vulcanization process of Plant 2. Flames and smoke were first observed under the pit of equipment No. 3115 on the 3100 line. The area showed severe damage and electrical abnormalities, and CCTV confirmed early smoke from the same location. The fire is presumed to have originated beneath the pit of No. 3115. 2 In the production building of a factory, sunflower oil was being refined by mixing oil, hexane, and activated carbon with a stirrer. Hexane vapors leaked and ignited from an unidentified ignition source, causing an explosion with fire. 3 During repair on the duct connected to a dust collector, grinder sparks ignited sponge soundproofing material on the wall. Scene investigation indicated initial ignition at the wall near the duct; burn patterns showed vertical spread to the second floor. The cause was determined as worker negligence. 4 Prior to an explosion and fire, an emergency call reported injuries due to steam leakage. Investigation found that unreacted photoinitiator was normally returned to a 5 ton reactor, but aluminum chloride catalyst was mistakenly added through a manhole, causing ejection. While workers evacuated, solvent vapors accumulated and later ignited from an unknown source, causing an explosion and fire. 5 A witness heard a beep and saw vapor leaking from a reactor lid. After donning a respirator, larger vapor was seen from the lid, condenser, and receiver tank area; droplets and chemicals fell from the reactor bottom. A brief initial explosion occurred, followed by a major explosion seconds later. 6 In a pharmaceutical production plant, an explosion and fire occurred during blending and drying of raw materials in the third floor mixing and granulation room. The fire spread and damaged a large portion of equipment and facilities. 7 A three story chemical plant producing BP F 95 experienced a fire presumed to have originated inside a centrifugal dehydrator while dehydrating BP F 95 mixed with toluene. The fire partially spread up to the third floor; no injuries were reported. 8 An explosion occurred due to a reaction between reactor contents and unidentified static electricity in a synthesis plant. The fire spread to nearby organic solvents, resulting in total destruction of the synthesis building. 9 In the drying room of Building D, an operator was loading wet powder cosmetic raw material into a dryer tray. Acetone vapors reached an explosive range and presumably ignited within the operator’s working radius, spreading to nearby combustibles. 10 No other heat source was identified at the origin except sparks from welding work. Burn patterns indicated rapid spread from stored flammable material (styrofoam). Based on witness statements and proximity to the welding site, welding sparks were presumed to have ignited the styrofoam.
By leveraging more than a decade of investigation reports (2013–2024), the analysis captures long-term and stable patterns of fire incidents within a relatively homogeneous industrial environment. This extensive temporal coverage mitigates sample-specific variability and enhances the generalizability of the derived semantic structures.
A total of 8,190 fire investigation reports were analyzed, from which 449 representative keywords were selected based on three filtering criteria: excluding 763 non-informative terms (e.g., numbers, symbols, general words, or names), consolidating 315 synonymous words into 165 representative entries, and retaining only those appearing at least twice across all reports.
@@ 4. Step 4
The Biterm Topic Model (BTM) estimates the probability of observing word pairs (biterms) across the entire corpus, rather than within individual documents. The resulting joint likelihood over all biterms $B$ is expressed as
@@
This formulation allows the model to capture global word co-occurrence patterns that help infer more stable topic–word associations. The conditional posterior for assigning a biterm to topic assignment $z$ is given by
@@
where $n_{z}$ is the number of biterms currently assigned to topic $z$, $n_{w|z}$ is the number of times word $w$ is assigned to topic $z$, and $\mathbf{z}_{-b}$ denotes topic assignments excluding the current biterm. Intuitively, this expression increases the probability of assigning a biterm to topics that already have high counts for its component words, while $\alpha$ and $\beta$ serve as smoothing hyperparameters controlling topic and word diversity.
To address convergence issues that may arise from direct Gibbs sampling, we employ collapsed Gibbs sampling, which integrates out $\bm{\theta}$ and $\bm{\phi}_{z}$ by exploiting conjugacy of the Dirichlet–Multinomial distributions [23]. After sufficient sampling iterations, the posterior estimates of topic–word distributions and topic proportions are:
@@
Here, $M$ denotes the total number of unique words in the corpus and $K$ represents the number of latent topics. Accordingly, the terms $M\beta$ and $K\alpha$ act as normalization constants arising from the Dirichlet–Multinomial conjugacy, where $\beta$ and $\alpha$ control the concentration of word and topic distributions, respectively. Larger values of $\beta$ or $\alpha$ lead to smoother, less peaked distributions, while smaller values encourage more concentrated topic–word associations.
Specifically, we construct a word–topic probability matrix $\mathbf{X}$ of dimension $N\times P$, where $N$ is the number of words and $P$ is the number of topics. To identify characteristic words, we compute two measures for each row of $\mathbf{X}$: (i) the coefficient of variation across topics, and (ii) the maximum probability. The coefficient of variation captures the relative dispersion of a word’s probabilities across topics, while the maximum probability identifies whether the word is strongly associated with at least one topic. Words with both high dispersion and high maximum probability are retained as representative terms, enabling more interpretable characterization of latent topics.
@@ 2.3 Clustering Topics via Latent Space Item Response Model
where $x_{i,p}$ denotes the probability of word $i$ belonging to topic $p$. The parameters $\bm{a}_{i}$ and $\bm{b}_{p}$ represent the intrinsic activity of word $i$ (i.e., how generally a word tends to appear across topics) and the popularity or overall strength of topic $p$, respectively. The Euclidean distance ${\bf v}_{i}-{\bf u}_{p}||$ measures how semantically close the word and topic are in the latent space, where smaller distances indicate stronger associations. Bayesian inference is employed to estimate the full parameter set $\bm{\Theta}=\{\bm{a},\bm{b},\mathbf{U},\mathbf{V}\}$ with appropriate priors, and parameters are sampled using Markov chain Monte Carlo (MCMC). As a result, we obtain latent positions of topics $\mathbf{v}_{i}$ in $\mathbb{R}^{d}$, forming the topic coordinate matrix ${\bf A}\in\mathbb{R}^{d\times P}$.
@@ 3.1 Topic Characterization through Words
As illustrated in Step 2 of Figure 1, the fire investigation texts were first analyzed through topic modeling to identify latent semantic structures. Table 2 presents the resulting thematic groups, providing an overview of how the documents are organized into distinct topics. In addition, Table 3 presents the distribution of words derived from the estimated topic–word probabilities ($\bm{\phi}_{w|z}$ obtained in the topic modeling stage. The table lists the top five terms with the highest posterior probabilities for each topic, highlighting representative keywords that characterize the semantic content of each domain.
@@
Table 3: Top five representative words for each extracted topic based on topic word probability distributions. These high probability words provide the basis for interpreting the thematic characteristics of the topics.
Topic Top 5 Words Topic 1 cleaning, cleaning room, wastewater treatment plant, agitator, machine room Topic 2 mixer, drum can, pallet, plastic, explosion Topic 3 electrical short circuit, electrical distribution, flame, circuit breaker, SWGR Topic 4 extractor hood, oven, laboratory, transformer, electricity Topic 5 heating wire, outdoor unit of air conditioner, thermal/acoustical insulation, air conditioner, rooftop Topic 6 dust collection equipment, duct, flame, plenty of, filter system Topic 7 storage, sesame dregs, residues, spontaneous combustion, storage room Topic 8 forklift, battery, electrical short circuit, distribution, engine compartment Topic 9 explosion, container, reactor, oil vapor, static electricity Topic 10 dust collection equipment, duct, flame, ventilation duct, plastic Topic 11 drying equipment, powder, silicon, base material, storage Topic 12 base material, mixing equipment, Banbury, thermal oil, flame Topic 13 welding, cooling tower, flame, sandwich panel, acetylene Topic 14 air compressor, motor, electrical short circuit, vulcanizer, solenoid valve Topic 15 waste materials, cigarette butts, waste, stacking, flame
@@ 3.2 Thematic Aggregation of Topics
Building upon the topic–word distributions, we further inferred interactions among words to explore higher-level thematic structures. Figure 2 presents the two-dimensional projection of the latent positions of the 449 words introduced in Section 2.1, with clusters identified using the $k$-means algorithm. The visualization highlights clusters by distinct colors, showing how semantically related words are grouped in close proximity. Based on model selection criteria, a total of 15 clusters were identified as valid. As shown in the Figure 2, words located near one another in the latent space tend to form coherent clusters, thereby capturing meaningful associations beyond the topic-level representation.
@@
Following Step 3 in Figure 1, the estimated latent positions of words were used to cluster semantically related terms based on their pairwise distances. Table 4 presents the top 10 representative keywords for each cluster, arranged according to their proximity to other words within the same cluster. This ordering enables a more precise characterization of the distinctive features of each cluster. While the identified clusters are not directly labeled as specific fire causes, their semantic coherence often reflects shared causal contexts. Accordingly, these clusters can be interpreted as data-driven representations of latent cause structures underlying fire incidents. Cluster 1 is anchored by the central term distillation column, which suggests the risk of explosion due to chemical leaks or failures in pressure control. Surrounding words such as exposure, solvent, heptane, and Silanes manufacturing reinforce the theme of hazardous material leakage leading to fires or explosions. Accordingly, Cluster 1 is labeled as Fire or explosion caused by hazardous material leakage.
Cluster 2 excludes the generic fire-related term ignition source and instead centers on polishing, which can produce ignition hazards when frictional heat contacts combustible materials. Neighboring words such as painting, floor, waste, and interior wall indicate that this cluster corresponds to Ignition from heat accumulation near combustible interior materials.
Cluster 3 is defined by the term incinerate, reflecting the inherent fire risks of incineration processes. Associated words such as waste wood, interior materials, heat, and urethane highlight ignition due to residual heat, while cutting machines and dust collection equipment suggest mechanical sources of smoldering. Thus, Cluster 3 is interpreted as Ignition due to residual heat post-work with combustibles.
Cluster 4 revolves around the term heat of reaction, but further interpretation requires contextualization with neighboring terms such as expired reagents, corn, and cooking oil. These materials, when awaiting disposal, pose risks of spontaneous combustion, particularly when combined with oxygen, heat waves, or rainwater. Hence, this cluster represents Spontaneous combustion from abandoned chemicals and oils.
Cluster 5 is structured around abnormal reaction, which links to numerous other terms and signals fires or explosions during abnormal chemical processing. Words such as pharmaceutical, film, oil, and epoxy point toward industrial settings where chemical instability can result in severe accidents. This cluster is labeled as Fires or explosions from abnormal reactions during chemical handling.
Cluster 6 includes terms like storage, cable, control box, and electrical circuit board, reflecting general electrical fires not tied to specific chemical processes. Its theme is summarized as Electrical fires in areas handling flammable materials.
Cluster 7 is characterized by coating, which signals risks associated with flammable paints. Surrounding terms such as corrosion, cutting oil, grinders, and presses suggest ignition by sparks or friction from industrial machinery. Accordingly, Cluster 7 is described as Ignition of flammable substances (e.g., machine oil) by friction heat or spark from machinery.
Cluster 8 is centered on vapor, a strong indicator of fire hazards in chemical plants. Terms such as nucleic acid, grease, and hazardous materials highlight risks arising from the ignition of volatile organic vapors, supporting the interpretation of Ignition of flammable vapor during organic solvent use.
Cluster 9 highlights the term drying room, referring to environments with elevated fire risks due to sustained high temperatures. Neighboring words, including cosmetic, ventilation fans, rotating, small amount, and decompose, suggest fire hazards associated with oil vapor and production processes. Thus, Cluster 9 is classified as Ignition from temperature rise in oil vapor areas.
Cluster 13 incorporates dilution, power outage, tracking, muller, vulcanizer, and melting fusion, suggesting interactions between hazardous materials, electrical ignition sources, and mechanical heat. This combination indicates Ignition in areas with accumulated combustible dust and oil vapors.
Cluster 14 includes terms such as micro, seat pad, deodorizing tower, absorbent pad, and gunnysack, which are not individually definitive. However, the presence of compost and sesame dregs highlights substances prone to spontaneous combustion through oxidation heat. Accordingly, this cluster represents Spontaneous combustion from improper oil residues storage or disposal.
Finally, Clusters 10, 11, 12, and 15 are each characterized by distinct keywords: Cluster 10 corresponds to Fire from ignition in accumulated combustibles, Cluster 11 to Fires from electrical factors in process equipment sites, Cluster 12 to Fires during machinery maintenance, and Cluster 15 to Ignition from sparks inside dust collection equipment with filters and debris. Importantly, these categories represent fire scenarios that extend beyond chemical plants to general factory environments.
@@
Table 4: Clusters of 10 representative keywords derived from interword correlations based on estimated latent positions.